OpenStudy (anonymous):

f(x) = x3 + 4x2 - x - 4 Part A: What are the factors of f(x)? Show your work. (3 points) Part B: What are the zeros of f(x)? Show your work. (2 points) Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). (5 points)

OpenStudy (anonymous):

someone please help!

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

Use synthetic division

OpenStudy (anonymous):

I don't know what that is

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

f(0)= -4 factors of -4 1,2,4 the long method use these values of x to find f(x) if f(x) = 0 then you have a factor when you get the first factor you can just use division to get a quadratic to factorize

OpenStudy (triciaal):

f(x) = x3 + 4x2 - x - 4 f(1) = 1 + 4 -1 - 4 = 0 therefore (x-1) is a factor

OpenStudy (anonymous):

was this for part a?

OpenStudy (triciaal):

yes part a means what are the x values for which f(x) = 0 1, -1, 4 f(x) = (x+1)(x-1)(x+4)

OpenStudy (triciaal):